12 or sister any 15 verse is your connected in series across a 9.0 v potential difference what is the current in the


1
Expert's answer
2021-05-24T14:10:43-0400

The given problem can be shown in below diagram:




R1 = 12 Ohms and R2 = 15 Ohms are in series,

Therefore, equivalent resistance R = R1 + R2

R = 12 + 15 = 27 Ohms

Voltage V = 9V

Therefore, the current I is given by:

"Current I = \frac{V}{R} = \frac{9}{27} = 0.33 Amp."


Hence the current in the circuit is I = 0.33 Amp.